### v3

# Dynamics of non-Hermitian Floquet Wannier-Stark system

Abstract
We study the dynamics of the non-Hermitian Floquet Wannier-Stark system in the framework of the tight-binding approximation, where the hopping strength is a periodic function of time with Floquet frequency $\mathcal{\omega}$. It is shown that the energy level of the instantaneous Hamiltonian is still equally spaced and independent of time $t$ and the Hermiticity of the hopping term. In the case of off resonance, the dynamics are still periodic, while the occupied energy levels spread out at the resonance, exhibiting $t^{z}$ behavior. Analytical analysis and numerical simulation show that the level-spreading dynamics for real and complex hopping strengths exhibit distinct behaviors and are well described by the dynamical exponents $z=1$ and $z=1/2$, respectively.
Keywords: Floquet Hamiltonian, non-Hermitian, resonance, anomalous diffusion, dynamical exponents, 2D simulator
1 Introduction
The Hermitian and static system is the main study object of traditional quantum mechanics, preserving the reality and conservation of energy. When these two constraints are violated, the properties of the material could be drastically modified, leading to new phases and phenomena [1, 2] without analogs in their static or Hermitian counterparts. Floquet engineering has facilitated the exploration of a diverse range of dynamic, topological, and transport phenomena across various physical contexts, including platforms involving ultracold atoms [3, 4, 5, 6]. More recently, the Floquet approach has been employed to engineer topological phases in non-Hermitian systems. Importantly, it was discovered that the interplay between time-periodic driving fields and the presence of gain, loss, or nonreciprocal effects can lead to the emergence of topological phases exclusive to non-Hermitian Floquet systems [7, 8, 9, 10, 11, 12, 13, 14, 15, 16] .
Motivated by the above investigations, the aim of the present paper is to examine how non-Hermiticity and periodic driving impact a Stark ladder system. Specifically, we study the dynamics of the non-Hermitian Floquet Wannier-Stark system in the framework of the tight-binding approximation. A conventional Wannier-Stark ladder consists of quantized equidistant energy levels, which separated by the Bloch frequency and emerge in a periodic Hermitian system under a linear potential [17, 18, 19]. Based on this framework, the Bloch oscillations in cold atoms within the optical lattice have also been theoretically investigated [20, 21, 22].
Recent works [23, 24, 25, 26] show that non-Hermitian hopping terms, including asymmtric and complex strengths, remain the reality of the Wannier-Stark ladder. On the other hand, it has proposed that a time-dependent hopping strength for a Bloch system can be realized by dynamically modulated waveguide arrays [27].
[…]
This paper is organized as follows. In Sec. 2, we introduce a non-Hermitian time-dependent tight-binding model supporting a Wannier-Stark ladder and present the analytical eigenstates of the instantaneous Hamiltonian. In Sec. 3, we focus on the dynamics of the static system, while the Floquet system at resonance in Sec. 4. Sec. 5 describes the application of the solutions to a specific initial state. In Sec. 6, a scheme for the simulator of our results is proposed in the framework of wavepacket dynamics in a $2$D square lattice. Finally, we summarize our results in Sec. 7.
Figure 1: (a) Schematic illustrations of the Hamiltonian in Eq. (1), which represents a tight-binding chain with a time dependence on the hopping strength between two adjacent sites. The on-site potential is tilted with slope $\omega_{0}$. It is shown that such a system supports a Stark-ladder with time-independent energy-level spacing. (b) Schematic illustrations of the Hamiltonian in Eq. (48), which is a tight-binding model of a square lattice with time-independent system parameters. The contributions of hopping strength and on-site potential can be used to simulate the dynamics of system (a).
2 Model and generalized Wannier-Stark ladder
We start with the tight-binding model with the Hamiltonian
$H=\kappa(t)\sum\limits_{n=-\infty}^{+\infty}(|n\rangle\langle n+1|+\text{ {H.c}})+\mathcal{\omega}_{0}\sum\limits_{n=-\infty}^{+\infty}n|n\rangle\langle n|,$ (1)
where $|n\rangle$ denotes a site state describing the Wannier state localized on the $n$th period of the potential. Here, $\kappa(t)$ is the time-dependent hopping strength, while $\mathcal{\omega}_{0}$ is the slope of the linear potential from a static force. Fig. 1(a) shows a schematic of the system. In our study, $\kappa(t)$ can be real and complex. In order to solve the time-dependent Hamiltonian, we start with the solution for the case where $\kappa$ is an arbitrary complex constant. We then discuss the time-dependent hopping strength $\kappa(t)$ at the end of the Section 3. The non-Hermitian Hamiltonians with complex hoppings have been realized in the coupled waveguides [31], optical fibres [32] and electrical circuit resonators [33].
It is well known that when the hopping constant $\kappa$ is real, the Hamiltonian ${H}$ in momentum space, based on the Bloch basis representation, can be exactly diagonalized. Consequently, the energy levels of ${H}$ are equidistant, which supports the periodic dynamics [34, 35, 36]. This phenomenon has been observed in systems of semiconductor superlattices and ultracold atoms [37, 38, 39, 40]. Now, we extend the conclusion to the case with arbitrary $\kappa$ through an alternative approach. The eigenstate of ${\left|\psi_{m}\right\rangle}$ can always be written in the form
${\left|\psi_{m}\right\rangle=\sum_{n}c_{nm}|n\rangle,}$ (2)
satisfying the Schrödinger equation
${H\left|\psi_{m}\right\rangle=E_{m}\left|\psi_{m}\right\rangle.}$ (3)
The eigenenergy $E_{m}$ can be obtained by the equation
$E_{m}={\kappa}\frac{c_{n+1,m}+c_{n-1,m}}{c_{nm}}+n\omega_{0},$ (4)
which accords with the recurrence relation of the Bessel function $J_{n}$ of argument $-2\kappa/\omega_{0}$, i.e.,
$m\omega_{0}={\kappa}\frac{J_{n-m+1}(-\frac{2\kappa}{\omega_{0}})+J_{n-m-1}(-\frac{2\kappa}{\omega_{0}})}{J_{n-m}(-\frac{2\kappa}{\omega_{0}})}+n\omega_{0}.$ (5)
Then, we have
$E_{m}=m\omega_{0},\ {c_{nm}=J_{n-m}(-\frac{2\kappa}{\omega_{0}}).}$ (6)
We can see that the eigenenergy ${m\mathcal{\omega}_{0}}$ is independent of $\kappa$, regardless of whether it is real or complex. This is crucial for the main conclusion of this work. Although the profile of the eigenstate
$\psi_{m}\left(n\right)=\langle n|\psi_{m}\rangle=J_{n-m}(-\frac{2\kappa}{\omega_{0}}),$ (7)
in real space is $\kappa$ dependent, and it is always localized. The inverse participation ratio (IPR) is a simple way to quantify the localization of a given state. For spatially extended states, the value of the IPR approaches zero when the system is sufficiently large, whereas it is finite for localized states regardless of the system size. The IPR of the eigenstate $\left|\psi_{m}\right\rangle$ is
${\ \text{IPR}_{m}=\frac{\sum\limits_{n}\left|\langle n|\psi_{m}\rangle\right|^{4}}{(\sum\limits_{n}\left|\langle n|\psi_{m}\rangle\right|^{2})^{2}}=\frac{\sum_{n}|J_{n-m}(-\frac{2\kappa}{\omega_{0}})|^{4}}{(\sum_{n}|J_{n-m}(-\frac{2\kappa}{\omega_{0}})|^{2})^{2}}.}$ (8)
[…]
Figure 2: Plots of the real and imaginary parts of energy levels for the system at different sizes. The parameters of the system are $\kappa=1+1i$ and $\omega_{0}=1$. One can observe that the real, equally spaced energy levels emerge from the center of the spectrum as the system size increases.
In parallel, without loss of generality, we have
$\left|\varphi_{m}\right\rangle=\sum_{n}J_{n-m}(-\frac{2\kappa^{\ast}}{\omega_{0}})|n\rangle,$ (9)
for the equation
$H^{{\dagger}}\left|\varphi_{m}\right\rangle=m\omega_{0}\left|\varphi_{m}\right\rangle,$ (10)
which establishes the biorthonormal set $\left\{|\varphi_{m}\rangle,|\psi_{n}\rangle\right\}$, satisfying
$\langle\varphi_{m}\left|\psi_{n}\right\rangle=\sum_{n^{\prime}}J_{n^{\prime}-m}^{\ast}(-\frac{2\kappa^{\ast}}{\omega_{0}})J_{n^{\prime}-n}(-\frac{2\kappa}{\omega_{0}})=\delta_{mn},$ (11)
since the eigenlevels are independent of $\kappa$, without coalescence of eigenvectors. This is an important basis for the following investigations.
The obtained solution to the non-Hermitian Hamiltonian obeys the well-established theory of non-Hermitian quantum mechanics [41, 42, 43, 44, 45]. In the context of non-Hermitian systems with $H\neq H^{{\dagger}}$, biorthonormal complete eigenstates refer to a set of eigenstates that are biorthonormal and form a complete basis for the Hilbert space of the system. The eigenstates of a non-Hermitian Hamiltonian can be complex and may not be orthogonal to each other. However, if a non-Hermitian Hamiltonian has a complete set of biorthonormal eigenstates, it means that: There exists a left eigenstate $\langle\varphi_{n}|$ and a right eigenstate $|\psi_{n}\rangle$, satisfying
$H|\psi_{n}\rangle=E_{n}|\psi_{n}\rangle,H^{{\dagger}}|\varphi_{n}\rangle=E_{n}^{\ast}|\varphi_{n}\rangle.$ (12)
for each eigenvalue $E_{n}$ such that
$\langle\varphi_{n}|\psi_{m}\rangle=\delta_{nm},\sum_{n}|\psi_{n}\rangle\langle\varphi_{n}|=1.$ (13)
The set of right eigenstates $\left\{|\psi_{n}\rangle\right\}$ and the set of left eigenstates $\left\{|\varphi_{n}\rangle\right\}$ together form a basis for the Hilbert space, allowing any state in the space to be expanded in terms of these eigenstates.
The concept of biorthonormal complete eigenstates is particularly important in non-Hermitian quantum mechanics because it plays an analogous role, both conceptually and computationally, to those in Hermitian quantum systems. For example, when we calculate the time evolution of a state, we first express the initial state $|\psi(0)\rangle$ as a linear combination of the right eigenstates
$|\psi(0)\rangle=\sum_{n}|\psi_{n}\rangle\langle\varphi_{n}|\psi(0)\rangle$ (14)
by projecting the initial state onto the left eigenstates. The state of the system at any later time $t$ is given by
$|\psi(t)\rangle=\sum_{n}|\psi_{n}\rangle e^{-iE_{n}t}\langle\varphi_{n}|\psi(0)\rangle=U(t)|\psi(0)\rangle$ (15)
where $U(t)$ is time evolution operator and will be obtained in the following for the present Hamiltonian in Eq. (1).
Figure 3: Plots of $P_{n}(t)/P(t)$ (see the main text) and $P(t)$ defined in Eqs. (19 ) and (20), obtained by numerical diagonalization for several typical values of system parameters with $\mathcal{\omega}_{0}=1$. Here, (a1), (b1), and (c1) correspond to the cases with $\omega=0$, $0.01$, and $0.1$, respectively, and $\kappa_{0}=1$. (a2), (b2), and (c2) correspond to the cases with $\omega=0$, $0.01$ , and $0.1$, respectively, and $\kappa_{0}=i$. (a3), (b3),and (c3) correspond to the cases with $\omega=0$ , $0.01$, and $0.1$, respectively, and $\kappa_{0}=e^{i\pi/4}$. We can see that the patterns in (a2, b2, c2) and (a3, b3, c3) are similar. This inidcates that the imaginary part of hopping plays an important role in the shapes of evolved states. According to the analytical analysis, ${P(t)}$ is periodic with the period depending on $\omega$ and $\omega_{0}$ When $\omega$ approaches $\omega_{0}$ , the period becomes infinite. The non-periodic behavior observed in (c1, c2, c3) arises because the scale plotted in the figure is smaller than the period.
Figure 4: Plots of $P_{n}(t)/P(t)$ (see the main text) and $\text{In}P(t)$ defined in Eqs. (19 ) and (20), obtained by numerical diagonalization for several typical values of system parameters with $\mathcal{\omega}_{0}=1$. Here, (a1), (b1), and (c1) correspond to the cases with $\omega=0.6$, $1$, and $1.5$, respectively, and $\kappa_{0}=1$. (a2), (b2), and (c2) correspond to the cases with $\omega=0.6$, $1$ , and $1.5$, respectively, and $\kappa_{0}=i$. (a3), (b3),and (c3) correspond to the cases with $\omega=0.6$, $1$, and $1.5$, respectively, and $\kappa_{0}=e^{i\pi/4}$.
3 Bloch oscillation in a non-Hermitian system
The dynamics of the system are governed by the propagator
$U_{mn}\left(t\right)=\langle m|e^{-iHt}|n\rangle.$ (16)
Employing the biorthonormal set $\left\{|\varphi_{m}\rangle,|\psi_{n}\rangle\right\}$, we have
$U_{mn}\left(t\right)$ $=$ $\sum_{l}\langle m|\psi_{l}\rangle e^{-iE_{l}t}\langle\varphi_{l}|n\rangle$ (17)
$=$ $J_{m-n}(-\frac{4\kappa}{\omega_{0}}\sin\frac{\omega_{0}t}{2})e^{-i(m-n)\frac{\omega_{0}t}{2}}i^{m-n}e^{-in\omega_{0}t},$
$=$ $i^{m-n}e^{-i(m+n)\frac{\omega_{0}t}{2}}J_{m-n}(-\frac{4\kappa}{\omega_{0}}\sin\frac{\omega_{0}t}{2}),$
which is a periodic function with period $2\pi/\omega_{0}$. For a given initial state
$\left|\Phi(0)\right\rangle=\sum_{l}c_{l}|l\rangle$ (18)
the probability distribution in real space for the evolved state $\left|\Phi(t)\right\rangle$ is
$P_{m}(t)=\left|\langle m\left|\Phi(t)\right\rangle\right|^{2}=\left|\sum_{l}c_{l}U_{ml}\left(t\right)\right|^{2},$ (19)
which is periodic. We obtain the conclusion that the total probability
$P(t)=\sum_{n}P_{n}(t),$ (20)
is still periodic, regardless of whether $\kappa$ is real or complex. In practice, the measurement of $P_{n}(t)$ has been reported in many non-Hermitian experiments [46, 47, 48, 49].
Up until now, we have been considering $\kappa(t)=\kappa$ as time-independent; now, we explore what happens when $\kappa(t)$ is time dependent. According to the adiabatic theorem, the expression of $U_{mn}\left(t\right)$ still holds true for slowly varying $\kappa(t)$ during a short period of time. Considering the case with $\kappa(t)=\kappa_{0}\cos(\omega t)$, where $\kappa_{0}$ can be a complex number and the frequency of $\kappa$ is small, $\mathcal{\omega}\ll\mathcal{\omega}_{0}$. It can be predicted that $\left|\Phi(t)\right\rangle$ and $P(t)$ are still approximately periodic even for the quasiadiabatic process. However, this approach could be invalid for the diabatic process.
To verify and demonstrate the above analysis, numerical simulations are performed to investigate the dynamic behavior of dynamic processes with different values of $\mathcal{\omega}/\mathcal{\omega}_{0}$. We compute the temporal evolution of a site state, $\left|\Phi(0)\right\rangle=|0\rangle$ by using a uniform mesh in the time discretization for the time-dependent Hamiltonian $H(t)$. The evolved state is
$\left|\Phi(t)\right\rangle=\mathcal{T}\exp\left[-i\int_{0}^{t}H\left(t\right)\mathrm{d}t\right]\left|\Phi(0)\right\rangle,$ (21)
where $\mathcal{T}$ is the time-order operator. Figs. 3 and 4 are the plots of $P_{n}(t)/P(t)$ and $P(t)$ for several typical values of $\omega/\omega_{0}$. For the non-Hermitian systems, the difference between the maximum and minimum of $P(t)$ is very large. Thus the profile of the evolved state cannot be properly presented by the plot of $P_{n}(t)$. To present a complete and visible profile, we plot the rescaled $P_{n}$ by $P_{n}(t)/P(t)$. When the total probability $P(t)$ is periodic in time, the periodicity of the evolved state can be confirmed by the profile of the probability distribution.
The numerical results indicate that the Bloch oscillation exists for small $\mathcal{\omega}/\mathcal{\omega}_{0}$. For non-Hermitian cases, the profile of the evolved state does not change substantially, but the probability is periodic. From Fig. 4 we can see that the evolution of the profile of the evolved state is still periodic when $\omega\mathbf{\neq}\omega_{0}$, but diffuses outward at $\omega=\omega_{0}$. This indicates that the dynamics are peculiar at resonance, which is the focus of the paper.
4 Floquet-resonance dynamics
Now, we investigate the dynamics of the system with time-dependent $\kappa(t)$. As previously mentioned, we consider periodically modulated $\kappa(t)=\kappa_{0}\cos(\omega t)$, where $\kappa_{0}$ can be a complex number. The solution of the Schrödinger equation
$i\frac{\partial}{\partial t}|\Psi\left(t\right)\rangle=H(t)|\Psi(t)\rangle$ (22)
can be written in the form
$|\Psi(t)\rangle=\sum_{m}a_{m}(t)e^{-im\mathcal{\omega}_{0}t}|\psi_{m}\rangle,$ (23)
where the coefficient $a_{n}$ obeys the following equation:
$\frac{\partial a_{m}}{\partial t}=-\sum_{n}a_{n}e^{i\left(m-n\right)\omega_{0}t}\left\langle\varphi_{m}\right|\frac{\partial}{\partial t}|\psi_{n}\rangle.$ (24)
Submitting the expression of $|\varphi_{m}(t)\rangle$ and $|\psi_{n}(t)\rangle$, we have
$\frac{\partial a_{n}}{\partial t}$ $=$ $\frac{1}{\omega_{0}}\frac{\partial\kappa}{\partial t}(a_{n-1}e^{i\mathcal{\omega}_{0}t}-a_{n+1}e^{-i\mathcal{\omega}_{0}t})$ (25)
$=$ $-\frac{{\kappa_{0}}\omega}{\omega_{0}}{\sin(\omega t)}(a_{n-1}e^{i\mathcal{\omega}_{0}t}-a_{n+1}e^{-i\mathcal{\omega}_{0}t}),$
which is the starting point for the following discussions.
Taking the rotating-wave approximation (RWA) under the condition $\left|\mathcal{\omega}-\mathcal{\omega}_{0}\right|\ll\mathcal{\omega}_{0}$, we have
$i\frac{\partial a_{n}}{\partial t}=\frac{\kappa_{0}\omega}{2\omega_{0}}[a_{n-1}e^{i(\mathcal{\omega}_{0}-\omega)t}+a_{n+1}e^{i(\omega-\mathcal{\ \omega}_{0})t}].$ (26)
We focus on the dynamics of the system at resonance $\mathcal{\omega}=\mathcal{\omega}_{0}$, in which the equation becomes
$i\frac{\partial a_{n}}{\partial t}=\frac{\kappa_{0}}{2}(a_{n-1}+a_{n+1}).$ (27)
The Schrödinger equation is used for a uniform tight-binding chain.
lt clearly shows that the eigenvalues of the equivalent Hamiltonian are proportional to the hopping constant $\kappa_{0}$, and then determines the behavior of the amplitude of the evolved state which can be conservative, damping, or explosive, respectively. The real and imaginary parts of $\kappa_{0}$ will affect the real and imaginary nature of the energy levels, which in turn will affect whether the probability is conserved and the evolution of the wave function. It can be anticipated that the presence of an imaginary part in $\kappa_{0}$ will significantly influence the dynamical evolution of the system.
5 Dynamical exponents
As shown above, the resonant dynamics are essentially governed by the simple equivalent Hamiltonian
[…]
where $|n\rangle$ denotes the eigenstates of the instantaneous Hamiltonian. The dynamics of $H_{\mathrm{eq}}$ were investigated in Ref. [25], which indicates that the reality of $\kappa_{0}$ plays an important role in the dynamics. The Hamiltonian $H_{\mathrm{eq}}$ can be diagonalized in the form
$H_{\mathrm{eq}}=\sum\limits_{k\in(-\pi,\pi)}\varepsilon_{k}|k\rangle\langle k|,$ (29)
with
$|k\rangle=\frac{1}{\sqrt{2\pi}}\sum\limits_{l}e^{ikl}|l\rangle,$ (30)
and the spectrum
$\varepsilon_{k}=\kappa_{0}\cos k.$ (31)
The time evolution of state
$\left|\Psi(t)\right\rangle=\sum\limits_{l}a_{l}(t)|l\rangle=\sum\limits_{l}a_{l}(0)\sum_{n}i^{n-l}J_{n-l}(-\kappa_{0}t)|n\rangle,$ (32)
can be obtained for any given $\left\{a_{l}(0)\right\}$. Intuitively, an initial state $|0\rangle$ (or $a_{l}(0)=\delta_{0l}$) could spread out and never return whether $\kappa_{0}$ is real or complex. In the following, we focus on two cases with $\kappa_{0}=1$ and $\kappa_{0}=i$.
Figure 5: The same plots of $P_{n}(t)/P(t)$ and $\text{In}P(t)$ as those in Fig. 3 are shown for several typical values of system parameters but at a resonance $\mathcal{\omega}_{0}=\omega=1$. Here, (a), (b), and (c) correspond to the cases with $\kappa_{0}=1$, $i$, and $e^{i\pi/4}$, respectively. The middle panels correspond to the probability distributions at moments $t=3T$, $6T$, and $9T$ , respectively. The right panels are the logarithmic plots of total probabilities as functions of time. These results agree with our analytical analysis in the main text. We can see that the patterns in (b) and (c) are almost identical. This inidcates that the imaginary part of hopping plays the important role in the energy-level spreading.
(i) $\kappa_{0}=1$; in this case, the straightforward derivation yields
$a_{n}(t)=i^{n}J_{n}(-t),$ (33)
which ensures the probability distribution of the energy levels
$\mathcal{P}_{n}(t)=\left|a_{n}(t)\right|^{2}=\left|J_{n}(-t)\right|^{2},$ (34)
and
$\mathcal{P}=\sum\limits_{n}\mathcal{P}_{n}(t)=1.$ (35)
The feature of the Bessel function indicates that there exists a maximum at the edge of $\mathcal{P}_{n}(t)$, which can be regarded as the wave front of the occupied-energy-level spreading. The location of such a wave front can then be determined by the equation
${\frac{\partial\mathcal{P}_{n_{c}}(t)}{\partial t}=0.}$ (36)
Based on the relations of Bessel functions
$J_{n_{c}-1}(x)-J_{n_{c}+1}(x)=2J_{n_{c}}^{\prime}(x),\,\,\,J_{n_{c}+1}\left(x\right)+J_{n_{c}-1}\left(x\right)=2\left|n_{c}/t\right|J_{n_{c}}\left(x\right),$ (37)
we have
${2J_{n_{c}+1}(-t)=2|n_{c}/t|J_{n_{c}}(-t).}$ (38)
For large $n$ and $t$, we have ${J_{n_{c}+1}(-t)\approx J_{n_{c}}(-t)}$, which results in
${|n_{c}|\approx t.}$ (39)
Then, we conclude that the dynamical exponent of occupied-energy-level spreading is $z=1$.
(ii) $\kappa_{0}=i$; in parallel, the corresponding amplitude is the Bessel function with imaginary argument
$a_{2n}(t)=(-1)^{n}J_{2n}(-it)=\frac{(-1)^{n}}{2\pi}\int_{-\pi}^{\pi}\cos(2nk)e^{-t\sin k}\mathrm{d}k,$ (40)
and
$a_{2n+1}(t)=i(-1)^{n}J_{2n+1}(-it)=\frac{(-1)^{n+1}}{2\pi i}\int_{-\pi}^{\pi}\sin[(2n+1)k]e^{-t\sin k}\mathrm{d}k.$ (41)
On a long time scale, the main contribution to the integral comes from the integrand around $k=-\pi/2$. Then, we have
$a_{2n}(t)\approx e^{-\frac{(2n)^{2}}{2t}}\frac{1}{2\pi}\int_{-\pi}^{\pi}e^{-t\sin k}\mathrm{d}k,$ (42)
and
$a_{2n+1}(t)\approx e^{-\frac{(2n+1)^{2}}{2t}}\frac{1}{2\pi i}\int_{-\pi}^{\pi}e^{-t\sin k}\mathrm{d}k,$ (43)
with small $\varepsilon_{k}$, which results in the probability distribution of the energy levels
$\mathcal{P}_{n}(t)=|a_{n}(t)|^{2}\approx\sqrt{\frac{1}{\pi t}}\mathcal{P}e^{-\frac{n^{2}}{t}},$ (44)
with the total probability
$\mathcal{P}=\sum_{n}\mathcal{P}_{n}=\frac{1}{2\pi}\int_{-\pi}^{\pi}e^{-2t\sin k}\mathrm{d}k.$ (45)
Obviously, the profile of $\mathcal{P}_{n}$ is a Gaussian type, and the total probability grows exponentially as $e^{2t}$ approaches the large $t$ limit. In this situation, the dynamics of occupied-energy-level spreading can be characterized by the full width at half maximum,
$\mathcal{P}_{n_{c}}(t)=\frac{1}{2}\mathcal{P}_{0}(t),$ (46)
which results in
$|n_{c}|\approx\sqrt{t}.$ (47)
Then, we conclude that the dynamical exponent of occupied-energy-level spreading is $z=1/2$.
The appendix Appendix presents the dynamical exponent for $\kappa_{0}$ as any complex number. When it is real, $z=1$; when it is not real, $z=1/2$. The non-Hermitian case differs from the Hermitian case primarily due to the imaginary part of $\kappa_{0}$. To verify and demonstrate the above analysis, numerical simulations are performed to investigate the dynamic behavior of dynamic processes at resonance with $\mathcal{\omega}=\mathcal{\omega}_{0}$. For the same initial state as shown in Fig. 3, the plots of $P_{n}(t)$ and $P(t)$ for several typical values of $\kappa_{0}$  are presented in Fig. 5. These numerical results agree with our above analysis in two aspects: (i) At resonance, the dynamics are no longer periodic; (ii) The level spreading exhibits distinct behaviors for Hermitian and non-Hermitian systems.
6 2D simulator via wavepacket dynamics
In this section, we will show how the time evolution of a wavepacket in an engineered time-independent 2D square lattice can simulate our results for Floquet dynamics in a 1D time-dependent lattice. In experiments, single-particle hopping dynamics can be simulated by discretized spatial light transport in an engineered 2D square lattice of evanescently coupled optical waveguides [50]. A 2D lattice can be fabricated by coupled waveguides, by which the temporal evolution of the single-particle probability distribution in the 2D lattice can be visualized by the spatial propagation of the light intensity.
We consider a single-particle Hamiltonian on a square lattice
$H_{\mathrm{2D}}$ $=$ $\kappa_{0}\sum_{n,m=-\infty}^{\infty}\cos\left(qm\right)|n,m\rangle\langle n+1,m|-J\sum_{n,m=-\infty}^{\infty}|n,m\rangle\langle n,m+1|+\mathrm{h.c.}$ (48)
$+\omega_{0}\sum_{n,m=-\infty}^{\infty}n|n,m\rangle\langle n,m|,$
which is an array of coupled 1D Bloch systems with $m$-dependent hopping strength $\kappa_{0}\cos\left(qm\right)$. Fig. 1(b) shows a schematic of the system. According to the above analysis, the system is governed by Bloch dynamics with frequency $\omega_{0}$ for vanishing $J$, regardless of whether $\kappa_{0}$ is real or complex.
Figure 6: For numerical simulations of the dynamics, the trajectories of wavepackets evolving in the in a 2D lattice with four typical configurations were obtained. The traces of the wavepacket $P(n,m)$ for the time evolutions of the four cases are shown in (a1), (b1), (c1), and (d1). The probability distributions $p(n,m,t)$ for the time evolutions of four representative cases at $t=0$, $t=2\pi/J$, $t=4\pi/J$, $t=6\pi/J$ and $t=8\pi/J\,$are shown in (a2), (b2), (c2), and (d2), respectively. The initial states and system parameters are given in the main text. The size of the system is $60\times 30$. The energy scale of the Hamiltonian is taken as $J=1$. The probability distributions on the square lattice exhibit distinct patterns corresponding to the dynamic behaviors of Bloch breathing, oscillation, diffusion and superdiffusion.
Now, we start to investigate the dynamics in the presence of nonzero $J$. The $k$-space representation of the Schrödinger equation
$i\frac{\partial}{\partial t}|\psi\rangle=H_{\mathrm{2D}}|\psi\rangle,$ (49)
can be written as
$i\frac{\partial\psi}{\partial t}$ $=$ $\kappa_{0}\cos k_{x}[\psi\left(k_{x},k_{y}+q,t\right)+\psi\left(k_{x},k_{y}-q,t\right)]$ (50)
$+i\omega_{0}\frac{\partial\psi}{\partial k_{x}}-2J\cos k_{y}\psi\left(k_{x},k_{y},t\right),$
by taking the Fourier transformation
$|k_{x},k_{y}\rangle=\frac{1}{2\pi}\sum_{n,m=-\infty}^{\infty}e^{i\left(k_{x}n+k_{y}m\right)}|n,m\rangle.$ (51)
For the case with $q=0$, the solution $\psi\left(k_{x},k_{y},t\right)$ reduces to
$\psi\left(k_{x},k_{y},t\right)=f\left(k_{x},t\right)e^{it2J\cos k_{y}},$ (52)
which resembles that in Ref. [51], where $f\left(k_{x},t\right)$ obeys the following equation:
$i\frac{\partial f}{\partial t}=(2\kappa_{0}\cos k_{x}+i\omega_{0}\frac{\partial}{\partial k_{x}})f\left(k_{x},t\right).$ (53)
In the case with $k_{y}\approx\pi/2$, we have $\cos k_{y}\approx k_{y}-\pi/2$. One can construct a solution by the superposition of states $|k_{x},k_{y}\rangle$ with $k_{y}\approx\pi/2$ in the form [52]
$\left|g\left(k_{x},t\right)\right\rangle$ $=$ $(\frac{2\alpha^{2}}{\pi})^{1/4}\int_{-\pi}^{\pi}e^{-\alpha^{2}\left(k_{y}-\pi/2\right)^{2}}e^{it2J\cos k_{y}}f\left(k_{x},t\right)\mathrm{d}k_{y}|k_{x},k_{y}\rangle$ (54)
$\approx$ $(\frac{1}{2\pi\alpha^{2}})^{1/4}\sum_{m}e^{im\pi/2}e^{-(1/4)(m-2Jt)^{2}/\alpha^{2}}f\left(k_{x},t\right)|k_{x},m\rangle,$
which is a traveling Gaussian wavepacket in the $y$ direction. It is presumed that such a solution still holds true by taking $\kappa_{0}\rightarrow\kappa_{0}\cos\left(qm\right)$ when $\kappa_{0}\cos\left(qm\right)$ varies slowly during the extension of the wavepacket in the y direction due to the locality of the solution. In this sense, the traveling wavepacket experiences an adiabatic process, after which the independence of the dynamics in both directions is maintained. Obviously, a small value of $q$ should be needed. On the other hand, the value of $\omega_{0}$ must match $q$ when the system is employed to simulate the dynamics at the resonance. To this end, a small value of $|\kappa_{0}|$ should be taken to maintain the Stark ladder.
Actually, under the conditions $|\kappa_{0}|\ll J$ and $k_{y}\approx\pi/2$, such a 2D system is equivalent to a coupled waveguide array due to the linear dispersion in the $y$ direction, in which the tunneling rate between neighboring waveguides is $m$ dependent. The speed of the propagating wavepacket along $m$ is $2J$. To simulate the dynamics along $x$, one can simply take $q=\omega/\left(2J\right)$, regarding $m$ as the temporal coordinate.
To simulate the time-dependent evolution of a 1D system (1) in a 2D lattice system (48), we design specific initial states
$\psi(n,m,0)=x(n)y(m),$ (55)
where $x(n)$ is taken as the initial state of the 1D system and $y(m)=({4\pi^{2}})^{-1/4}e^{im\pi/2}e^{-m^{2}/4}$ represents a Gaussian wave packet with momentum ${\pi}/{2}$ in the y-direction, serving as an alternative to the time axis. At first, we verify this scheme by considering the Bloch oscillation in Hermitian system. (i) To simulate Bloch breathing, $x(n)$ is taken as $\delta_{n,n_{0}}$. The system parameters are $\kappa_{0}=J$, $\omega=0$ and $\omega_{0}=0.5J$. (ii) To simulate Bloch oscillation of a wavepacket, $x(n)$ is taken as $e^{-{(n-n_{0})^{2}}/4}$. The system parameters are $\kappa_{0}=J$, $\omega=0$, and $\omega_{0}=0.5J$. Second, we consider the simulations for the dynamics, which are presented in Fig. 4(b1) and (b2). (iii) To simulate superdiffusion, $x(n)$ is taken as $\delta_{n,n_{0}}$. The system parameters are $\kappa_{0}=0.25J$ and $\omega=\omega_{0}=0.5J$. (iv) To simulate diffusion, $x(n)$ is taken as $\delta_{n,n_{0}}$. The system parameters are $\kappa_{0}=0.25Ji$ and $\omega=\omega_{0}=0.5J$. For greater clarity, the specific parameters are shown in Tab. 1. The evolved state has the form
$\left|\psi\left(n,m,t\right)\right\rangle=e^{-iH_{\mathrm{2D}}t}\left|\psi\left(n,m,0\right)\right\rangle,$ (56)
which can be computed by exact diagonalization.
Table 1: List of the parameters of the four driven Hamiltonians and initial states, where $y(m)=({4\pi^{2}})^{-1/4}e^{im\pi/2}e^{-m^{2}/4}$ represents a Gaussian wave packet with momentum ${\pi}/{2}$ in the y-direction, moving at a constant velocity $v=2J$ in the y-direction. Therefore, the distance the wave packet travels in the y-direction represents time.
       $\kappa_{0}$        $\omega$        $\omega_{0}$        Initial state
[…]
The probability distribution at position $(n,m)$ at time $t$ is defined by the probabilities
$p(n,m,t)=\left|\langle n,m,\left|\psi\left(n,m,t\right)\right\rangle\right|^{2}.$ (57)
To visualize the profile of the dynamics, we define the function
$P(n,m)=\sum_{j=1}p(n,m,j\tau),$ (58)
to record the trace of the wavepacket. Here, we use $j\tau$ to denote the discretized time coordinate. It is expected that the distribution of $P(n,m)$  exhibits patterns similar to those in Fig. 5. The numerical results presented in Fig. 6 show that the probability distributions on the square lattice exhibit distinct patterns corresponding to the dynamic behaviors of Bloch breathing, oscillation, superdiffusion and diffusion.
7 Conclusion
In summary, we simultaneously examine the effects of non-Hermiticity and periodic driving on a Stark ladder system. The most fascinating and important feature of such systems is that the Stark ladder can be maintained regardless of the non-Hermiticity of the hopping strength, and the energy level spacing depends only on its norm. This allows us to solve the problem in an analytical manner. The Bloch oscillation is broken when the Floquet frequency is resonant with the energy level spacing, and the occupied energy levels spread out. Furthermore, analytic analysis and numerical simulation show that the level-spreading dynamics for real and complex hopping strengths exhibit distinct behaviors of superdiffusion and diffusion, respectively. In addition, we propose a scheme to simulate the results by wavepacket dynamics in a 2D square lattice. This can be experimentally realized in the coupled array of waveguides. These findings may help us deepen our understanding of the interplay between non-Hermiticity and periodic driving impacts on the dynamics of the system.
Acknowledgements
Funding information
This work was supported by the National Natural Science Foundation of China (under Grant No. 12374461).
Appendix
In this appendix Appendix, we derive the dynamical exponent for the case in which $\kappa_{0}$ is an arbitrary complex number. When $\kappa_{0}$ is a real number, i.e., $\text{Im}(\kappa_{0})=0$, using the same Bessel function estimation method as in the main text, one can obtain the dynamical exponent of occupied-energy-level spreading $z=1$, and the total probability $\mathcal{P}$ remains constant at 1.
When $\text{Im}(\kappa_{0})\neq 0$, we have
$a_{2n}(t)=(-1)^{n}J_{2n}(-\kappa_{0}t)=\frac{(-1)^{n}}{2\pi}\int_{-\pi}^{\pi}\cos(2nk)e^{i\kappa_{0}t\sin k}\mathrm{d}k,$ (59)
and
$a_{2n+1}(t)=i(-1)^{n}J_{2n+1}(-\kappa_{0}t)=\frac{(-1)^{n+1}}{2\pi i}\int_{-\pi}^{\pi}\sin[(2n+1)k]e^{i\kappa_{0}t\sin k}\mathrm{d}k.$ (60)
On a long time scale, the main contribution to the integral comes from the integrand around $k=-\pi/2$(or $\pi/2$, depends on the sign of $\text{Im}(\kappa_{0})$). Then, we have
$a_{2n}(t)$$\approx$ ${e^{-\frac{(2n)^{2}}{2i\kappa_{0}t}}\frac{1}{2\pi}\int_{-\pi}^{\pi}e^{i\kappa_{0}t\sin k}\mathrm{d}k,\,\text{Im}(\kappa_{0})<0}$ (61)
$\approx$ ${e^{\frac{(2n)^{2}}{2i\kappa_{0}t}}\frac{1}{2\pi}\int_{-\pi}^{\pi}e^{i\kappa_{0}t\sin k}\mathrm{d}k,\,\text{Im}(\kappa_{0})>0}$ (62)
and
$a_{2n+1}(t)$$\approx$ ${e^{-\frac{(2n+1)^{2}}{2i\kappa_{0}t}}\frac{i}{2\pi}\int_{-\pi}^{\pi}e^{i\kappa_{0}t\sin k}\mathrm{d}k,\,\text{Im}(\kappa_{0})<0}$ (63)
$\approx$ ${e^{\frac{(2n+1)^{2}}{2i\kappa_{0}t}}\frac{1}{2\pi i}\int_{-\pi}^{\pi}e^{i\kappa_{0}t\sin k}\mathrm{d}k,\,\text{Im}(\kappa_{0})>0}$ (64)
with small $\varepsilon_{k}$, which results in the probability distribution of the energy levels
$\mathcal{P}_{n}(t)=|a_{n}(t)|^{2}\approx\sqrt{\frac{|\text{Im}(\kappa_{0})|}{\pi|\kappa_{0}|^{2}t}}\mathcal{P}e^{-\frac{|\text{Im}(\kappa_{0})|n^{2}}{|\kappa_{0}|^{2}t}},$ (65)
with the total probability
$\mathcal{P}=\sum_{n}\mathcal{P}_{n}=\frac{1}{2\pi}\int_{-\pi}^{\pi}e^{-2\text{Im}(\kappa_{0})t\sin k}\mathrm{d}k.$ (66)
Obviously, the profile of $\mathcal{P}_{n}$ is a Gaussian type, and the total probability grows exponentially as $e^{2|\text{Im}(\kappa_{0})|t}$ approaches the large $t$ limit. In this situation, the dynamics of occupied-energy-level spreading can be characterized by the full width at half maximum,
$\mathcal{P}_{n_{c}}(t)=\frac{1}{2}\mathcal{P}_{0}(t),$ (67)
which results in
$|n_{c}|\approx\sqrt{t}.$ (68)
Then, we conclude that the dynamical exponent of occupied-energy-level spreading is $z=1/2$.